The difference between two numbers is 16. Four times the smaller number is equal to 8 more than the larger. What are the numbers?
Answer by macston(5194)

x=smaller number: y=larger number
y-x=16
y=x+16
4x=y+8 Substitute for y
4x=(x+16)+8 Subtract x from each side.
3x=24 Divide each side by 3.
x=8 ANSWER 1: The smaller number is 8.
y=x+16=8+16=24 ANSWER 2: The larger number is 24.
CHECK:
4x=y+8
4(8)=24+8
32=32